OpenStudy (photon336):

Find an equation for the tangent lines to the hyperbola x^2-y^2 = 16 that pass through (2,-2).

OpenStudy (photon336):

I got the first part down. I'm just not sure how to find the equation of the line that goes through (2,-2) and touches the hyperbola \[x^{2}+y^{2} = 16\] \[\frac{ d }{ dx }x^{2}+\frac{ d }{ dx }y^{2} = 0 \] \[2x+2y~\frac{ dy }{ dx } = 0 \] \[\frac{ dy }{ dx } = \frac{ -2x }{ -2y }\] \[\frac{ dy }{ dx } = \frac{ x }{ y }\]

OpenStudy (freckles):

how did you get the extra negative factor ?

OpenStudy (photon336):

err... wait.. sorry it was x^{2}-y^{2} = 16

OpenStudy (freckles):

ok lol the other thing was a circle anyways :p that makes more sense

OpenStudy (photon336):

yeah, Like i don't know what to do after this.. it's saying something like find an equation for the tangent line that also passes through the point (2,-2)

OpenStudy (freckles):

so say the line goes through (a,f(a)) and (2,-2) where (a,f(a)) is the point of tangent in question

OpenStudy (photon336):

yep following

OpenStudy (freckles):

\[a^2-(f(a))^2=16 \\ -(f(a))^2=16-a^2 \\ (f(a))^2=a^2-16 \\ f(a)= \pm \sqrt{a^2-16} \] so we have two assumptions to make about f(a) \[\text{ assume } f(a) >0 \\ \frac{dy}{dx}|_{(x,y)=(a,\sqrt{a^2-16})} =\frac{a}{\sqrt{a^2-16}} \\ \text{ but we also can find the slope of the line algebraically } \\ \text{ that is using } \frac{\sqrt{a^2-16}-(-2)}{a-2}\]

OpenStudy (freckles):

so we have an equation to solve

OpenStudy (freckles):

we will go back later and make the assumption y is negative

OpenStudy (freckles):

i mean f(a)

OpenStudy (photon336):

really ridiculous question: on the second line what happened to the negative sign? Other question is this: okay so you're saying that two points a and f(a) are on the circle. and that we insert those points into the equation for a circle and then isolate to say find one of the points?

OpenStudy (freckles):

hyperbola :p also we are trying to find the point of tangency in question a point of tangency will happen on the relation/function so yes it is on the hyperbola we are using the calculus definition of slope and using the algebraic definition of slope to solve for a

OpenStudy (freckles):

and one sec about the negative sign question

OpenStudy (freckles):

are you talking about going from 2nd to 3rd or from 1st to 2nd?

OpenStudy (freckles):

1st to 2nd I subtract something on both sides 2nd to 3rd I multiplied both sides by -1

OpenStudy (photon336):

-(f(a))^{2} = 16-a^{2}

OpenStudy (freckles):

I'm going to draw a graph of what we are doing I only drew one possible tangent line that contained (2,-2) there looks like there is 3 more